# On single-frequency asymptotics for

### 19 Mar 2026 v1

@@ Abstract
We calculate all harmonic states and analyse their stability. The calculations rely on the Bloch–Feynman gyroscopic representation of von Neumann equation for the density matrix. The asymptotics follow by application of the averaging theory of the Bogolyubov type.
@@ 1 Introduction
The Maxwell–Bloch equations (MBE) were introduced by Lamb [24] for the semiclassical description of the laser action [1, 6, 17, 31, 32, 33, 34]. The equations are the Galerkin approximation of the Maxwell–Schrödinger system [5, 16, 18, 20, 21, 26, 29] Our main goal is construction of solutions with single-frequency asymptotics. The asymptotics seem to correspond to the laser coherent radiation, which remains a key mystery of laser action since its invention around 1960. The damped driven MBE for mixed states read
@@
In Appendix A, we prove the a priori bound for solutions $X(t)=(A(t),B(t),\rho(t))$ to (1.1):
$|X(t)|\leq D_{1}|X(0)|e^{-\Gamma t}+D_{2},\quad t>0;\qquad D_{1},D_{2},\Gamma>0,$ (1.5)
which implies the well-posedness of the MBE, and existence of a compact global attractor.
@@
Our main results are the following asymptotics for solutions $X(t)=(M(t),S(t))$ to equations (1.1) (in the representation (1.8)) with a fixed quotient $p/\gamma=r$. Everywhere below we consider the case $p>0$ only since the extension to $p<0$ is obvious.
Theorem 1.1.
@@
The peculiarity of the asymptotics (1.11)– (1.12) is that they hold only for a fixed quotient $p/\gamma=r$, and do not hold without this restriction. The asymptotics (1.11) for $S(t)$ reads
@@
The asymptotics (1.11) follows for solutions with every initial state $(M(0),S(0))\in Z^{r}$ by the averaging theory of Bogolyubov type [30, Theorem 4.3.6]. To prove asymptotics (1.12), we combine the theorem with the stability of the branch $Z^{r}_{+}$. Our calculations show that for all harmonic states $({\bf M},{\bf S})\in Z^{r}_{+}$, the component ${\bf M}=-{\bf A}^{e}$. This is why the limiting amplitude $\bf M$$_{*}$ in (1.12) does not depend on $r>0$ and on $X(0)\in D^{r}_{d}$.
@@ 2 The Bloch-Feynman “gyroscopic” representation
Here we represent von Neumann equation from MBE as a “gyroscopic equation” (2.6) applying the Rabi, Bloch, Feynman & al approach [6, 7, 14, 15, 28]. Hermitian density matrices $\rho$ satisfying (1.2), admit the representation
@@
Hence, using the Pauli matrices, we obtain (1.8):
@@
The nonnegativity of $\rho$ is equivalent to
$\det\rho=1-S_{1}^{2}-S_{2}^{2}-S_{3}^{2}\geq 0.$ (2.3)
[…]
Using completeness of the Pauli matrices, we can expand the “Hamiltonian operator” (1.2) as
@@ 4 The averaging
which implies nonexistence of stationary states with nonzero Maxwell field for the system (4.1), and also the decay ${\bf M}(t)=\displaystyle{\bf M}(0)e^{-\frac{\gamma}{2}t}$ that agrees with (A.1) and (A.3). In contrast, the resonance equation (4.4) includes the interaction term which can prevent the decay of the Maxwell amplitude, that is expected physically.
@@ 5 Stationary states for the averaged dynamics (harmonic states)
It is important that the stationary equations depend on $r$ but do not depend on $p$. The first equation of (5.1) gives
@@ 8.1 KBM vector field
Let us denote $v_{r}=(f_{r},g_{r})$ the vector field (3.10) of the system (3.9). It is easy to check that in the case of almost periodic pumping $A^{e}(t)$ and each $r\!\neq\!0$, for any bounded region $C{\cal D}\subset{\mathbb C}\times{\mathbb C}$, we have
@@
Moreover, $v_{r}$ is the Lipschitz continuous vector field in any bounded region $C{\cal D}\subset{\mathbb C}\times{\mathbb C}$. Hence, $v_{r}$ is a KBM (Krylov–Bogolyubov–Mitropolsky) vector field in any bounded region $\cal D$} according to [30, Definition 4.2.4].
@@ 8.2 The asymptotics in the interaction picture
Here we prove Theorem 1.1.
@@
The initial state $({\mathbb M}(0),{\mathbb S}(0))=(M(0),S(0))$ by (3.3). Hence, $({\mathbb M}(0),{\mathbb S}(0))=({\bf M},{\bf S})$. However, $({\bf M},{\bf S})$ is a stationary state for the averaged system (1.9), or equivalently, to (4.7). Therefore, Theorem 4.3.6 of [30] implies that
$\quad\,\,\,\,\,\max_{t\in[0,p^{-1}]}|({\mathbb M}(t),{\mathbb S}(t))-({\bf M},{\bf S})|={\cal O}({p^{1/2}}),\qquad p\to 0,\quad{p}/\gamma=r$ (8.4)
since both conditions 1 and 2 of the theorem hold in our case since
a) $v_{r}$ is a KBM-vector field in any bounded region $C{\cal D}\subset{\mathbb C}\times{\mathbb C}$ with the order function (8.2);
b) each solution is bounded due to the conservation (2.8) and the a priori estimate (A.1).
Hence, (8.3) follows.
@@ Appendix A The a priori bounds and well-posedness
Let $\Omega\neq 0$ and $A^{e}(t)\in C[0,\infty)$. Then for solutions to (1.1), the following bounds hold:
$|A(t)|+|B(t)|\leq d_{1}(|A(0)|+|B(0)|)e^{-\Gamma t}+d_{2},\,\,\,\,t>0;\quad d_{1},d_{2},\Gamma>0.$ (A.1)
Proof.
The first two equations of (1.1) are equivalent to the ODE
$\ddot{A}(t)=-\Omega^{2}A(t)-\gamma\dot{A}(t)+cj(t).$ (A.2)
The solutions are given by
$A(t)=A_{1}e^{\lambda_{1}t}+A_{2}e^{\lambda_{2}t}+c\int_{0}^{t}\frac{e^{\lambda_{1}(t-s)}-e^{\lambda_{2}(t-s)}}{\lambda_{1}-\lambda_{2}}j(s)ds,\qquad\lambda_{1,2}=-\frac{\gamma}{2}\pm\sqrt{\frac{\gamma^{2}}{4}-\Omega^{2}}.$ (A.3)
Now (A.1) follows since $\Omega\neq 0$, while $|j(t)|=2\varkappa|\rho_{21}(t)|\leq 2\varkappa$ by (1.1) and (1.2), and $B(t)=\dot{A}(t)$. ∎
Remark A.2.
The bound (1.5) implies that there exists the compact global attractor for the MBE which is the minimal subset ${\cal A}\subset{\mathbb X}\,\,$ such that
$X(t)\to{\cal A},\qquad t\to\infty$
for any solution $X(t)$ to (1.1). This follows by known methods [4, 12, 22, 35].

### 23 Mar 2026 v2

@@ Abstract
We calculate all harmonic states and analyse their stability. The calculations rely on the Bloch–Feynman gyroscopic representation of von Neumann equation for the density matrix. The asymptotics follow by application of the averaging theory of the Bogolyubov type. The key role in the application of the averaging theory is played by a special a priori estimate.
@@ 1 Introduction
The Maxwell–Bloch equations (MBE) were introduced by Lamb [25] for the semiclassical description of the laser action [1, 6, 17, 31, 32, 33, 34]. The equations are the Galerkin approximation of the Maxwell–Schrödinger system [5, 16, 19, 21, 22, 27, 29] Our main goal is construction of solutions with single-frequency asymptotics. The asymptotics seem to correspond to the laser coherent radiation, which remains a key mystery of laser action since its discovery around 1960. The damped driven MBE for mixed states read
@@
In Appendix A, we prove the a priori bound for solutions $X(t)=(A(t),B(t),\rho(t))$ to (1.1) with $|p|/\gamma=r$ satisfy the following a priori bounds:
$|X(t)|\leq D_{r}(|X(0)|),\qquad t\geq 0,$ (1.5)
which is proved in Appendix A. The bound implies the well-posedness of the MBE.
@@
Our main results are the following asymptotics for solutions $X(t)=(M(t),S(t))$ to equations (1.1) in the representation (1.8), with a fixed quotient $p/\gamma=r$. Everywhere below we consider the case $p>0$ only since the extension to $p<0$ is obvious.
Theorem $1.$
@@
Remark 1.2.
Without the constraint $p/\gamma=r$, the a priori bounds (1.5) and the asymptotics (1.11)–(1.12) do not hold. For example, the bounds must be different for small and big $|p|$ for a given dissipation $\gamma>0$. The asymptotics do not hold without this restriction since the limiting amplitudes ${\bf M}^{r}$ depend on $r$.
The asymptotics (1.11) for $S(t)$ reads
@@
The asymptotics (1.11) follows for solutions with every initial state $(M(0),S(0))\in Z^{r}$ by the averaging theory of Bogolyubov type [30, Theorem 4.3.6]. To prove asymptotics (1.12), we combine the theorem with the stability of the branch $Z^{r}_{+}$. Our calculations show that for all harmonic states $({\bf M},{\bf S})\in Z^{r}_{+}$, the component ${\bf M}=-{\bf A}^{e}$. This is why the limiting amplitude $\bf M$$_{*}$ in (1.12) does not depend on $r>0$ and on $X(0)\in D^{r}_{d}$. The key role in the application of the averaging theory is played by the special a priori estimate (1.5).
@@ 2 The Bloch-Feynman “gyroscopic” representation
Here we represent von Neumann equation from MBE as a “gyroscopic equation” (2.5) applying the Bloch, Feynman & al approach [6, 7, 14, 15]. Hermitian density matrices $\rho$ with $\mathop{\rm tr\,}\nolimits\rho=1$ admit the representation
@@
Hence, using the Pauli matrices and the nonnegativity $\det\rho=1-S_{1}^{2}-S_{2}^{2}-S_{3}^{2}\geq 0$, we obtain (1.8):
@@
In particular, we can expand the “Hamiltonian operator” (1.2) as
@@ 4 The averaging
which implies nonexistence of stationary states with nonzero Maxwell field for the system (4.1), and also the decay ${\bf M}(t)=\displaystyle{\bf M}(0)e^{-\frac{\gamma}{2}t}$. In contrast, the resonance equation (4.4) includes the interaction term which can prevent the decay of the Maxwell amplitude, that is expected physically.
@@ 5 Stationary states for the averaged dynamics (harmonic states)
It is important that the stationary equations depend on $r$ but do not depend on pp. The first equation of (5.1) gives
@@ 8.1 KBM vector field
Let us denote $v_{r}=(f_{r},g_{r})$ the vector field (3.10) of the system (3.9). It is easy to check that in the case of almost periodic pumping $A^{e}(t)$ and each $r\!\neq\!0$, for any bounded region ${\cal D}\subset{\mathbb X}$, we have
@@
Moreover, $v_{r}$ is the Lipschitz continuous vector field in any bounded region ${\cal D}\subset{\mathbb X}$. Hence, $v_{r}$ is a KBM (Krylov–Bogolyubov–Mitropolsky) vector field in any bounded region $\cal D$} according to [30, Definition 4.2.4].
@@ 8.2 The asymptotics in the interaction picture
Here we prove Theorem $1.$
@@
The initial state $({\mathbb M}(0),{\mathbb S}(0))=(M(0),S(0))$ by (3.3). Hence, $({\mathbb M}(0),{\mathbb S}(0))=({\bf M},{\bf S})$. However, $({\bf M},{\bf S})$ is a stationary state for the averaged system (1.9), or equivalently, to (4.7). Therefore, Theorem 4.3.6 of [30] implies (8.3). Indeed, both conditions 1 and 2 of the theorem hold in our case since
a) $v_{r}$ is a KBM-vector field in any bounded region ${\cal D}\subset{\mathbb X}$ with the order function (8.2);
b) solutions $({\mathbb M}(t),{\mathbb S}(t))$ of the system (3.5) with ${p}/\gamma=r$ and small $p>0$ are uniformly bounded by the a priori estimate (A.4).
Hence, (8.3) is proved.
@@ Appendix A The a priori bounds and well-posedness
There exists a Lyapunov function $V(A,B)$ such that
$a_{1}[A^{2}+B^{2}]\leq V(A,B)\leq a_{1}[A^{2}+B^{2}]\quad{\rm where}\quad a_{1},a_{2}>0,$ (A.1)
and for solutions to (1.1), the function $V(t)=V(A(t),B(t))$ satisfies the inequality
$\dot{V}(t)\leq-\gamma bV(t)+d\frac{p^{2}}{\gamma},\qquad t>0;\quad b,d>0.$ (A.2)
Corollary A.2.
Solving the inequality (A.2), we obtain:
$V(t)\leq V(0)+\frac{d}{b}r^{2},\qquad t\geq 0.$ (A.3)
Hence, for solutions to (1.1) with $p/\gamma=r$, the following bounds hold:
$A^{2}(t)+B^{2}(t)\leq D_{r}(A(0),B(0)),\qquad t\geq 0.$ (A.4)
Now (1.5) follows.